Home‎ > ‎Crew 136‎ > ‎

228-CommandersReport

Commander Report
02/28/2014
Corentin Liber

Today was our last full day at the MDRS.  We decided to spend it out of SIM conditions to have the
opportunity to check the equipment and make sure everything is functional for the next crew.

We went on EVA in the morning for a last check on the streambed study area project while the other
crew members worked on the finalization of their reports.  We spent most of the afternoon working on
the last processing needed for some projects and writing our reports.  Some maintenance also took place.

We went to pick up Chuck in Hanksville around 3.30 pm and brought him back to the Hab with us.  He
will stay until tomorrow.

-----------------------------

CommanderCheckIn Report
02/28/2014
Thibaud Herriau

Crew Physical Health:
All are well.
Eva Depart Time:
8:50 am
Eva Return Time:
10:30 am
Narrative Of Field Mission Results:
Data collected about the hardness and nature of the soil in the streambed.

Hab Maint:
Multiple failures in the greenhab.
Rep Sched:
None
Tomorrow Plan:
Cleaning MDRS
Welcoming crew137 and briefing each member on the different roles
(Engineer, HSO, Greenhab officer etc.) and on MDRS in general.
Inventory:
None
Support Req:
None
Miscelleneous:
None